# Properties of observable mixed inertial/gravito-inertial modes in $\gamma$ Doradus stars

### v1

@@ 3.3 Evolution of the width and central spin parameter of the dips
Figure 7: Period spacing in the co-rotating frame as a function of the spin parameter for the modes of the series ($\ell=1$, $m=-1$) at different rotation rates using the 1z model. The period spacing is normalised by the buoyancy radius $\Pi_{0}$. The pink line shows the traditional approximation of rotation. The square dots indicate the modes that are within the range of typically observed radial orders $7030<n<$ 70. The dashed line corresponds to the frequency of the resonant inertial mode $\ell_{i}=3$, $m=-1$ calculated analytically with the uniform-density model.
@@
Figure 8: Same as Fig. 7 for the model 2m. The low-amplitude oscillation in $\Delta P$ are due to a glitch generated by a small numerical discontinuity in its Brunt-Väisälä frequency profile (see Miglio et al., 2008).
Figure 9: Same as Fig. 7 for the model 3t.
@@
Figure 11: Width of the dips $\sigma$ as a function of the rotation rate for the series ($\ell=1$, $m=-1$), for the three stellar models described in Table 1. Error bars are too small to be visible with the scale used.
@@
We show in Figs. 7, 8 and 9 the evolution of the first dip of the series ($\ell=1$, $m=-1$) with the star rotation rate for the 1z, 2m and 3t model respectively. The square dots indicate modes within an observable range of radial orders. Indeed, according to Li et al. (2020), the observed modes of the ($\ell=1$, $m=-1$) series are typically of radial order $7030<n<$ 70. Analogous figures for the other two series ($\ell=2$ $m=-2$) and ($\ell=1$, $m=0$) are shown in Appendix C.
@@
We fit the dips using Eq. (9) to determine their width $\sigma$ and central spin parameter $s_{c}$. Figures 10 and 11 shows the evolution of $s_{c}$ and $\sigma$ with the star rotation for the three stellar models. The label ”core-only model” refers to a model that will be later discussed in Sect. 4.2, related oscillation modes are shown in Appendix B. We see on Fig. 10 that $s_{c}$ varies with the rotation rate but becomes less sensitive to it as the star evolves. It generally increases with the rotation rate, with the exception of the 3t model between $\Omega=0.1\Omega_{\mathrm{K}}$ and $\Omega=0.15\Omega_{\mathrm{K}}$. We also see that $s_{c}$ decreases as the star evolves. The width $\sigma$ follows the same trend: it increases with the rotation rate but becomes smaller and less sensitive to the rotation rate as the star evolves (see Fig. 11). In summary, the rotation rate of the star tend to make the dips occur at higher $s$ and widen them, whereas the dips become narrower and occur at a lower $s$ as the star ages.
@@
We notice that outside of the dip and for the three models we consider, the curve remains further and further from the TAR as the rotation rate grows. This behaviour has already been observed in previous studies (e.g., Ouazzani et al., 2017).
@@ 4.1 Empirical expression of $\sigma$
According to the model of Tokuno & Takata (2022), $\sigma$ evolves linearly with the parameter $\epsilon=\Omega/N_{0}$ where $N_{0}$ is the jump of the Brunt-Väisälä frequency at the interface between the convective core and the radiative region. We can test whether $\sigma$ is proportional to $\Omega$ by plotting its evolution with $\Omega$ alongside the theoretical linear evolution predicted in Tokuno & Takata (2022). Figure 12 shows the evolution of $\sigma/\sigma_{\mathrm{ref}}$ with the rotation rate where $\sigma_{\mathrm{ref}}$ is the width computed at $\Omega=\Omega_{\mathrm{ref}}=0.2\Omega_{\mathrm{K}}$. This shows that $\sigma$ is indeed proportional to the rotation rate of the star, with the exception of the 3t model for which we can see a deviation that is significant compared to the errors of $\sigma$. The origin of this particular behaviour remains unclear at the moment and numerical effects cannot be excluded. Analogous figures for the series ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) are shown in Appendix C.
Figure 12: Evolution of $\sigma/\sigma_{\mathrm{ref}}$ with the rotation rate, for the dip of the series ($\ell=1$, $m=-1$) for three different evolutionary stages (1z, 2m, 3t). The associated error bars are too small to be visible. The reference value $\sigma_{\mathrm{ref}}$ is the width for $\Omega=\Omega_{\mathrm{ref}}=0.2\Omega_{\mathrm{K}}$. If, as predicted by Tokuno & Takata (2022), $\sigma\propto\Omega$, $\sigma/\sigma_{\mathrm{ref}}=\Omega/\Omega_{\mathrm{ref}}=5\Omega$ (the dashed line).
Then, assuming a linear relation $\sigma=\beta\epsilon$, we perform a linear regression to determine the slope $\beta$. We use the least-square method. Table 2 gathers the values of $\beta$ obtained for each dip and stellar model. While $\beta$ is expected to depend on the core structure (Tokuno & Takata, 2022), we find that it varies by less than $15$% for the three stellar models considered. For the dip of the ($\ell=1$, $m=-1$) series, we show in Fig. 13 the result of the fit obtained by considering the three stellar models together. When directly fitting with relation $\sigma=\beta\epsilon$, the 1z-model $\sigma$ values weight more in the determination of the free parameter $\beta$ because, due to the smaller $N_{0}$ of the 1z model, they extend over a larger $\epsilon$ range. To counter that, we actually fit the relation $\sigma N_{0}=\beta\Omega$. Analogous figures for the dips of the ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) series are shown in Appendix C.
Table 2: Values of $\beta$ obtained by fitting the relation $\sigma N_{0}=\beta\Omega$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$), for the models 1z, 2m and 3t separately and together.
@@
Figure 13: Width $\sigma$ as a function of $\epsilon$ for the series ($\ell=1$, $m=-1$) for 3 different evolutionary stages (1z, 2m, 3t). The purple dashed line shows the fit presented in Table 2 for the three models together and the purple area shows the related error. The x-axis is logarithmic for visualisation purposes.
Using our empirical relation, the mean value of $\beta$ could thus be used to estimate $N_{0}$ from $\sigma$ via
$N_{0}^{\mathrm{calc}}=\frac{\Omega\beta}{\sigma}.$ (10)
@@ 4.2 Empirical expression of $s_{c}$
In order to understand the variations of $s_{c}$, we first computed the spin parameters $s_{*}$ of inertial modes of truncated versions of our models limited to their convective cores, as done in Ouazzani et al. (2020). It allows us to see the impact of varying density as well as the effect of the rotation rate on the frequency of the resonant inertial modes. They are calculated with the boundary condition $\xi_{r}=0$, $\xi_{r}$ being the radial displacement, and identified as the resonant modes using their number of latitudinal and radial nodes. The inertial modes and their spin parameter obtained for this core-only model are reported in Appendix B. As already observed in Ouazzani et al. (2020), the spin parameter of these modes remains independent of $\Omega$ as long as an acoustic term can be neglected in the governing equations (see Eq. (23)).
@@
We can then compare the central spin parameter of the dip $s_{c}$ obtained in the previous section with the spin parameter $s_{*}$ of the resonant inertial mode of the core-only model. We show the comparison for the series ($\ell=1$, $m=-1$) for the three stellar models in Fig. 10. The spin parameter of inertial modes obtained using the uniform-density model presented in Appendix A is also indicated. Analogous figures for the ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) series are shown in Appendix C. We notice that for a same inertial mode, the spin parameter decreases with the star evolution. This behaviour depends on the density stratification of the core, which is related to the stellar mass and the stellar evolution. We see that while the spin parameters of the eigenmodes of the uniform-density model and the core-only model remain constant with the rotation (with the exception of $s_{*}$ for the 3t model between $\Omega=0.1\Omega_{\mathrm{K}}$ and $\Omega=0.15\Omega_{\mathrm{K}}$), the central spin parameter of the dip shows a significant evolution with the rotation rate. For the three models, $s_{c}$ tends toward $s_{*}$ as the rotation rate is decreasing. Additionally, the differences between the full model and the core-only model are significantly decreasing as the star ages. It is consistent with the fact that the Brunt-Väisälä jump $N_{0}$ grows as the star evolves on the main sequence which makes the radial displacement $\xi_{r}$ at the core interface increases, making the core more isolated, thus closer to the core-only model. This also explains the decrease of the dip width with evolutionary stage as a less deformable interface weakens the coupling, which thus impacts less and less modes and leads to narrower dips.
@@
Looking for an empirical relationship between $s_{c}$ and $\epsilon$ usable on observation data, we noticed that, while $s_{\rm c}$ does not vary linearly with the rotation for all models (especially 3t), the difference $s_{c}-s_{*}$ appears to be proportional to $\Omega$ (figure not shown). We thus fit $s_{c}-s_{*}$ as a linear function of $\epsilon$: $s_{c}-s_{*}=\alpha\epsilon$. We estimated $\alpha$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$) and for the models 1z, 2m, and 3t separately. Table 3 provides the estimate of $\alpha$ for each case.
Table 3: Values of $\alpha$ obtained fitting the relation $s_{c}(\epsilon)-s_{*}(\epsilon)=\alpha\epsilon$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$), for the models 1z, 2m and 3t.
@@
We can see that $\alpha$ varies significantly (up to $\sim$30%) with the star model for a same series of modes.
However, we observed that the ratio $s_{c}/s_{*}$ plotted as a function of $\epsilon$ appears to almost gather on a same line for the three models (see Fig. 14). Since this quantity depends less on the model, we fit it as a linear function $s_{c}/s_{*}=A\epsilon+B$ for the three models, separately and together. We obtain values of B compatible with 1, this confirms that $s_{c}$ tends to $s_{*}$ when $\epsilon$ tends to 0. Therefore we redo fits by fixing $B=1$, letting only $A$ as a free parameter. As we did for the estimation of $\beta$, we fit $(s_{c}/s_{*}-1)N_{0}=A\Omega$ so all the models weight equally. The result of the fit for $A$ is shown in Table 4.
Table 4: Values of $A$ obtained by fitting the relation $(s_{c}/s_{*}-1)N_{0}=A\Omega$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$), for the models 1z, 2m and 3t, separately and together.
@@
The global fit of all models together is plotted in Fig. 14 for the series ($\ell=1$, $m=-1$). Analogous figures for the ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) series are shown in Appendix C. As $A$ does not vary very much with the stellar structure, we assume it to be constant in order to retrieve the frequency of the resonant inertial mode in the core, $s_{*}$, from the location of an observed dip through the relation
@@
Using Eq. (10), we can rewrite this relation
@@
As $s_{c}$ and $\sigma$ are obtained from the $\Delta P_{\mathrm{co}}$-$s$ relation in the co-rotating frame, the rotation rate $\Omega$ is needed. It can be determined by fitting the TAR outside the dip (e.g. Li et al., 2020).
Figure 14: Ratio $s_{c}/s_{*}$ as a function of $\Omega/N_{0}$ ($=\epsilon$) for the series ($\ell=1$, $m=-1$), for the 1z, 2m and 3t models. The purple dashed line shows the fit presented in Table 4 for the three models together and the purple area shows the related error. The x-axis is logarithmic for visualisation purposes.
4.3 Relative errors made on the estimation of $N_{0}$ and $s_{*}$

### v2

@@ 3.3 Evolution of the width and central spin parameter of the dips
Figure 7: Period spacing in the co-rotating frame as a function of the spin parameter for the modes of the series ($\ell=1$, $m=-1$) at different rotation rates using the 1z model. The period spacing is normalised by the buoyancy radius $\Pi_{0}$. The pink line shows the traditional approximation of rotation. The square dots indicate the modes that are within the range of typically observed radial orders $7030<n<$ 70. The dashed line corresponds to the frequency of the resonant inertial mode $\ell_{i}=3$, $m=-1$ calculated analytically with the uniform density model.
@@
Figure 8: Same as Fig. 7 but for the model 2m. The low-amplitude oscillations in $\Delta P$ are due to a glitch generated by a small numerical discontinuity in its Brunt-Väisälä frequency profile (see Miglio et al., 2008).
Figure 9: Same as Fig. 7 but for the model 3t.
Figure 10: Evolution of the centre of the dip $s_{\mathrm{c}}$ obtained using a full model (full line) and evolution of the spin parameter of the inertial mode obtained using a core-only model (dotted line). The dash-and-dot blue line shows the spin parameter of the inertial mode in the case of a uniform density core (analytical model). The light blue curves refer to the model 1z, orange curves to the model 2m, and red curves to the model 3t.
@@
Figure 11: Width of the dips $\sigma$ as a function of the rotation rate for the series ($\ell=1$, $m=-1$) for the three stellar models described in Table 1. The error bars are too small to be visible with the scale used.
@@
We show in Figs. 7, 8, and 9 the evolution of the first dip of the series ($\ell=1$, $m=-1$) with the star rotation rate for the 1z, 2m, and 3t model, respectively. The square dots indicate modes within an observable range of radial orders. Indeed, according to Li et al. (2020), the observed modes of the ($\ell=1$, $m=-1$) series are typically of a radial order $7030<n<$ 70. Analogous figures for the other two series, ($\ell=2$ $m=-2$) and ($\ell=1$, $m=0$), are shown in Appendix C.
@@
We fit the dips using Eq. (9) in order to determine their width $\sigma$ and central spin parameter $s_{c}$. Figures 10 and 11 show the evolution of $s_{c}$ and $\sigma$ with the star rotation for the three stellar models. The label ‘core-only model’ refers to a model that we discuss later in Sect. 4.2 (some oscillation modes calculated with this model are shown in Appendix B). One can see in Fig. 10 that $s_{c}$ varies with the rotation rate but becomes less sensitive to it as the star evolves. It generally increases with the rotation rate, with the exception of the 3t model between $\Omega=0.1\Omega_{\mathrm{K}}$ and $\Omega=0.15\Omega_{\mathrm{K}}$. We also observed that $s_{c}$ decreases as the star evolves. The width $\sigma$ follows the same trend: It increases with the rotation rate but becomes smaller and less sensitive to the rotation rate as the star evolves (see Fig. 11). In summary, the rotation rate of the star tends to make the dips occur at a higher $s$ and widens them, whereas the dips become narrower and occur at a lower $s$ as the star ages.
@@
We note that outside of the dip and for the three models we considered, the curve remains further and further from the TAR as the rotation rate grows. This behaviour has already been observed in previous studies (e.g. Ouazzani et al., 2017).
@@ 4.1 Empirical expression of $\sigma$
According to the model of Tokuno & Takata (2022), $\sigma$ evolves linearly with the parameter $\epsilon=\Omega/N_{0}$, where $N_{0}$ is the jump of the Brunt-Väisälä frequency at the interface between the convective core and the radiative region. We tested whether $\sigma$ is proportional to $\Omega$ by plotting its evolution with $\Omega$ alongside the theoretical linear evolution predicted in Tokuno & Takata (2022). Figure 12 shows the evolution of $\sigma/\sigma_{\mathrm{ref}}$ with the rotation rate, where $\sigma_{\mathrm{ref}}$ is the width computed at $\Omega=\Omega_{\mathrm{ref}}=0.2\Omega_{\mathrm{K}}$. This figure shows that $\sigma$ is indeed proportional to the rotation rate of the star, with the exception of the 3t model for which one can see a deviation that is significant compared to the errors of $\sigma$. The origin of this particular behaviour remains unclear at the moment, and numerical effects cannot be excluded. Analogous figures for the series ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) are shown in Appendix C.
Figure 12: Evolution of $\sigma/\sigma_{\mathrm{ref}}$ with the rotation rate for the dip of the series ($\ell=1$, $m=-1$) for three different evolutionary stages (1z, 2m, 3t). The associated error bars are too small to be visible. The reference value $\sigma_{\mathrm{ref}}$ is the width for $\Omega=\Omega_{\mathrm{ref}}=0.2\Omega_{\mathrm{K}}$. If, as predicted by Tokuno & Takata (2022), $\sigma\propto\Omega$, then $\sigma/\sigma_{\mathrm{ref}}=\Omega/\Omega_{\mathrm{ref}}=5\Omega$ (the dashed line).
Then, assuming a linear relation $\sigma=\beta\epsilon$, we performed a linear regression to determine the slope $\beta$. We used the least-squares method. Table 2 gathers the values of $\beta$ obtained for each dip and stellar model. While $\beta$ is expected to depend on the core structure (Tokuno & Takata, 2022), we find that it varies by less than $15$% for the three stellar models considered. For the dip of the ($\ell=1$, $m=-1$) series, we show in Fig. 13 the result of the fit obtained by considering the three stellar models together. When directly fitting with relation $\sigma=\beta\epsilon$, the 1z-model $\sigma$ values are weighted more in the determination of the free parameter $\beta$ because, due to the smaller $N_{0}$ of the 1z model, they extend over a larger $\epsilon$ range. To counter that, we actually fitted the relation $\sigma N_{0}=\beta\Omega$. Analogous figures for the dips of the ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) series are shown in Appendix C. Using our empirical relation, the mean value of $\beta$ could thus be used to estimate $N_{0}$ from $\sigma$ via
$N_{0}^{\mathrm{calc}}=\frac{\Omega\beta}{\sigma}.$ (10)
Table 2: Values of $\beta$.
@@
 [^2] Values of $\beta$ are obtained by fitting the relation $\sigma N_{0}=\beta\Omega$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$), and ($\ell=2$, $m=-2$) for the models 1z, 2m, and 3t separately and together.
Figure 13: Width $\sigma$ as a function of $\epsilon$ for the series ($\ell=1$, $m=-1$) for three different evolutionary stages (1z, 2m, 3t). The purple dashed line shows the fit presented in Table 2 for the three models together, and the purple area shows the related error. The x-axis is logarithmic for visualisation purposes.
@@ 4.2 Empirical expression of $s_{c}$
In order to understand the variations of $s_{c}$, we first computed the spin parameters $s_{*}$ of inertial modes of truncated versions of our models limited to their convective cores, as done in Ouazzani et al. (2020). It allowed us to see the impact of varying the density as well as the effect of the rotation rate on the frequency of the resonant inertial modes. The modes were calculated with the boundary condition $\xi_{r}=0$, with $\xi_{r}$ being the radial displacement, and identified as the resonant modes using their number of latitudinal and radial nodes. The inertial modes and their spin parameter obtained for this core-only model are reported in Appendix B. As already observed in Ouazzani et al. (2020), the spin parameter of these modes remains independent of $\Omega$ as long as an acoustic term can be neglected in the governing equations (see Eq. (23)).
@@
We could then compare the central spin parameter of the dip $s_{c}$ obtained in the previous section with the spin parameter $s_{*}$ of the resonant inertial mode of the core-only model. We show the comparison for the series ($\ell=1$, $m=-1$) for the three stellar models in Fig. 10. The spin parameter of the inertial modes obtained using the uniform density model presented in Appendix A is also indicated. Analogous figures for the ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) series are shown in Appendix C. We note that for the same inertial mode, the spin parameter decreases with the star evolution. This behaviour depends on the density stratification of the core, which is related to the stellar mass and the stellar evolution. We observed that while the spin parameters of the eigenmodes of the uniform-density model and the core-only model remain constant with the rotation (with the exception of $s_{*}$ for the 3t model between $\Omega=0.1\Omega_{\mathrm{K}}$ and $\Omega=0.15\Omega_{\mathrm{K}}$), the central spin parameter of the dip shows a significant evolution with the rotation rate. For the three models, $s_{c}$ tends towards $s_{*}$ as the rotation rate decreases. Additionally, the differences between the full model and the core-only model significantly decrease as the star ages. This is consistent with the fact that the Brunt-Väisälä jump $N_{0}$ grows as the star evolves on the main sequence, which makes the radial displacement $\xi_{r}$ at the core interface increase and thus makes the core more isolated and closer to the core-only model. This also explains the decrease of the dip width with evolutionary stage, as a less deformable interface weakens the coupling, which thus impacts fewer and fewer modes and leads to narrower dips.
@@
While looking for an empirical relationship between $s_{c}$ and $\epsilon$ usable on observation data, we noticed that though $s_{\rm c}$ does not vary linearly with the rotation for all models (especially 3t), the difference $s_{c}-s_{*}$ appears to be proportional to $\Omega$ (figure not shown). We thus fit $s_{c}-s_{*}$ as a linear function of $\epsilon$: $s_{c}-s_{*}=\alpha\epsilon$. We estimated $\alpha$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$), and ($\ell=2$, $m=-2$) and for the models 1z, 2m, and 3t separately. Table 3 provides the estimate of $\alpha$ for each case.
Table 3: Values of $\alpha$.
@@
 [^3] Values of $\alpha$ are obtained by fitting the relation $s_{c}(\epsilon)-s_{*}(\epsilon)=\alpha\epsilon$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$), and ($\ell=2$, $m=-2$) for the models 1z, 2m, and 3t.
One can see that $\alpha$ varies significantly (up to $\sim$30%) with the star model for a same series of modes. However, we observed that the ratio $s_{c}/s_{*}$ plotted as a function of $\epsilon$ appears to almost gather on a same line for the three models (see Fig. 14). Since this quantity depends less on the model, we fitted it as a linear function $s_{c}/s_{*}=A\epsilon+B$ for the three models, separately and together. We obtained values of B compatible with 1, and this confirms that $s_{c}$ tends to $s_{*}$ when $\epsilon$ tends to 0. Therefore, we redid fits by fixing $B=1$, letting only $A$ be a free parameter. As we did for the estimation of $\beta$, we fitted $(s_{c}/s_{*}-1)N_{0}=A\Omega$ so that all the models were weighted equally. The result of the fit for $A$ is shown in Table 4.
Table 4: Values of $A$.
@@
 [^4] Values of $A$ are obtained by fitting the relation $(s_{c}/s_{*}-1)N_{0}=A\Omega$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$), and ($\ell=2$, $m=-2$) for the models 1z, 2m, and 3t, separately and together.
@@
The global fit of all models together is plotted in Fig. 14 for the series ($\ell=1$, $m=-1$). Analogous figures for the ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) series are shown in Appendix C. As $A$ does not vary much with the stellar structure, we assumed it to be constant in order to retrieve the frequency of the resonant inertial mode in the core, $s_{*}$, from the location of an observed dip through the following relation:
@@
Using Eq. (10), we rewrote this relation as
@@
Since $s_{c}$ and $\sigma$ are obtained from the $\Delta P_{\mathrm{co}}$-$s$ relation in the co-rotating frame, the rotation rate $\Omega$ is needed. It can be determined by fitting the TAR outside the dip (e.g. Li et al., 2020).
Figure 14: Ratio $s_{c}/s_{*}$ as a function of $\Omega/N_{0}$ ($=\epsilon$) for the series ($\ell=1$, $m=-1$) for the 1z, 2m, and 3t models. The purple dashed line shows the fit presented in Table 4 for the three models together, and the purple area shows the related error. The x-axis is logarithmic for visualisation purposes.
4.3 Relative errors made on the estimations of $N_{0}$ and $s_{*}$